OpenStudy (anonymous):

PLEASE HELP ASAP !!!! 4x+6y=62 -6x+6y=-18

OpenStudy (anonymous):

Does it have options?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

A)(6,8) b(8,4) C(6,6) d(8,5)

OpenStudy (anonymous):

D (8,5) STOP YELLING AT ME! </3

OpenStudy (anonymous):

SORRY :(

OpenStudy (anonymous):

4x+6y=62 -6x+6y=-18 --------------- - 10x=80 x=8 4x+6y=62 4(8) + 6y=62 6y=62-32 6y=30 y=5